OpenStudy (anonymous):

help please! how do you solve this quadratic equation by completing the square: -x^2+6x+10=0

OpenStudy (nottim):

What adds to six and multiplies to 10.

OpenStudy (nottim):

Or is it the other way around?

hero (hero):

-x^2 + 6x + 10 = 0 -(x^2 - 6x - 10) = 0 x^2 - 6x - 10 = 0 x^2 - 6x = 10 x^2 - 6x + 9 = 10 + 9 (x - 3)^2 = 29 x - 3 = + or - sqrt(29) x = + or - sqrt(29) + 3

OpenStudy (anonymous):

this isn't that kind of problem. you are suppose to make it be a perfect square. I got to this... -x^2+6x+10=0 -10 -x^2+6x=-10 Now you have to force it to become a perfect square. soo... -x^2+6x+9=-10+9 sqrt -(x+3)^2= sqrt -1 -x+3= sqrt -1 -x=-3 + or - sqrt -1 then I am confused!

hero (hero):

^well, you were supposed to add 10 to both sides, then you would have gotten 10 + 9 = 29 instead of -1

hero (hero):

By the way, I did make it a perfect square. I know how to "complete the square" and that's what I did.

OpenStudy (anonymous):

nooo you have subtract -10 and then add nine because you had to subtract to cancel out the ten to make it zero on the left side so zero minus 10 is -10 and then you have to make the perfect square & you did. but you just messed up on the end part. I just don't understand how to simplify the ending. My teacher did not go over this well in class..

OpenStudy (anonymous):

But I was replying to @NotTim

OpenStudy (anonymous):

But thank you anyways, :)

hero (hero):

Trust me @apeymarie93, I know what I'm talking about.

hero (hero):

@apeymarie93

hero (hero):

You failed to understand my first two steps

OpenStudy (pokemon23):

LOL what a coincidence me and hero was doing this

hero (hero):

That would be the appropriate course of action that would ultimately lead to having to add 10 to both sides, rather than subtract it.

hero (hero):

Uh, no Pokie..this is complete the square

hero (hero):

@pokemon23

OpenStudy (anonymous):

Ohh I see! How would do this one then? x^2+2x-7 I got. x= -1 + or - 2 sqrt 2

OpenStudy (pokemon23):

@Hero THEIR MORE

hero (hero):

x^2 + 2x - 7 = 0 x^2 + 2x = 7 this is the same thing that I was explaining with pokie earlier....the leading term must be positive before proceeding...In this case, x^2 is positive, so yes, we can simply add 7 to both sides. if it were: -x^2 + 2x - 7 = 0 we would have to do this: -(x^2 -2x + 7) = 0 x^2 - 2x + 7 = 0 x^2 - 2x = -7 Do you get the difference now?

hero (hero):

I probably confused you. I should have just posted the solution first.

OpenStudy (anonymous):

I am completely confused...

hero (hero):

I know, and I apologize. I was just trying to help you understand the difference of approach when you have something like x^2 + 6x + 7 = 0 vs -x^2 + 6x + 7 = 0

hero (hero):

The first step is not the same with those two

OpenStudy (anonymous):

Oh so if you have -x^2 you have to factor out the negative?.....?

hero (hero):

precisely

OpenStudy (anonymous):

So like did i do the other one right? My teacher went over this in like 10 minutes Friday. I get like the all the steps now, but like.. I don't understand the very last step.

hero (hero):

which very last step? Why don't you post your full solution to this and then put an arrow next to the step you don't get?

OpenStudy (anonymous):

Like where you have to simplify the square root. Hold on one second.

OpenStudy (anonymous):

x^2+2x -7=0 x^2 +2x + 1=8 sqrt(x+1)^2 = sqrt 8 x= -1 + or - sqrt(8) How do you simplify it down?

hero (hero):

You don't

hero (hero):

That's the end of it

OpenStudy (anonymous):

But the answer in the back of my book says. x= -1 + or - 2 sqrt (2)

hero (hero):

I was just about to mention that...sqrt(8) = sqrt(4*2) = 2 sqrt(2) that's why I came back.

OpenStudy (anonymous):

See, that's what I don't understand..